# Heisenberg limit in phase measurements: the threshold detection approach

### v1

@@ Appendix A Threshold detection in the two-arm interferometer case
Let the states $|{\psi_{k_{1},k_{2}}}\rangle$, where $k_{1,2}=1,\dots$, complement $|{\psi_{0,0}}\rangle$ to the two-dimensional full orthonormal set:
@@
$\langle{k_{1},k_{2}}|\hat{\mathcal{V}}^{\dagger}\hat{Y}\hat{\mathcal{V}}|{l_{1},l_{2}}\rangle=\begin{cases}0\,,&k_{1}=k_{2}=l_{1}=l_{2}=0\,,\\ \langle{0,0}|\hat{\mathcal{Y}}_{2}|{l_{1},l_{2}}\rangle\,,&k_{1}=k_{2}=0\,,\ l_{1}+l_{2}>0\,,\\ \langle{k_{1},k_{2}}|\hat{\mathcal{Y}}_{2}^{\dagger}|{0,0}\rangle\,,&k_{1}+k_{2}>0\,,\ l_{1}=l_{2}=0\,,\\ \text{undefined}\,,&k_{1}+k_{2}>0\,,\ l_{1}+l_{2}>0\,,\end{cases}$ (102)
@@
$\hat{\mathcal{V}}^{\dagger}\hat{Y}\hat{\mathcal{V}}=\hat{\mathcal{Y}}_{2}+\hat{\mathcal{Y}}_{2}^{\dagger}=-2[\sqrt{2}\alpha\hat{p}_{2}e^{r}+(\hat{x}_{1}\hat{p}_{2}+\hat{x}_{2}\hat{p}_{1})\mathop{\rm sinh}\nolimits R_{+}].$ (105)
@@
Substituting Eqs. (64) into Eq. (81), we obtain:
$\hat{Y}_{\rm out}=2\bigl{\{}2\alpha^{2}e^{r-R}\bigl{(}\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\tfrac{1}{2}\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi\bigr{)}\hfill\\ +\sqrt{2}\alpha\bigl{[}\hat{x}_{1}e^{r}(\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi)-\hat{p}_{2}e^{-R}(\mathop{\rm cosh}\nolimits R_{+}\cos\phi+\mathop{\rm sinh}\nolimits R_{+}\cos 2\phi)\bigr{]}\\ +\bigl{[}\tfrac{1}{2}\bigl{(}(\hat{x}_{1}^{2}+\hat{x}_{2}^{2})e^{R_{+}}-(\hat{p}_{1}^{2}+\hat{p}_{2}^{2})e^{-R_{+}}\bigr{)}\sin 2\phi-(\hat{x}_{1}\hat{p}_{2}+\hat{x}_{2}\hat{p}_{1})\cos 2\phi\bigr{]}\mathop{\rm sinh}\nolimits R_{+}\bigr{\}}.$ (107)
Therefore,
$\hat{Y}_{\rm out}|{0,0}\rangle=2\bigl{\{}2\alpha^{2}e^{r-R}\bigl{(}\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\tfrac{1}{2}\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi\bigr{)}\\ +\alpha\bigl{[}\hat{a}_{1}^{\dagger}e^{r}(\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi)-i\hat{a}_{2}^{\dagger}e^{-R}(\mathop{\rm cosh}\nolimits R_{+}\cos\phi+\mathop{\rm sinh}\nolimits R_{+}\cos 2\phi)\bigr{]}\\ +\bigl{[}\bigl{(}\tfrac{1}{2}(\hat{a}_{1}^{\dagger}{}^{2}+\hat{a}_{2}^{\dagger}{}^{2})\mathop{\rm cosh}\nolimits R_{+}+\mathop{\rm sinh}\nolimits R_{+}\bigr{)}\sin 2\phi-i\hat{a}_{1}^{\dagger}\hat{a}_{2}^{\dagger}\cos 2\phi\bigr{]}\mathop{\rm sinh}\nolimits R_{+}\bigr{\}}|{0,0}\rangle,$ (108)
and
$\langle{\hat{Y}_{\rm out}}\rangle=4\alpha^{2}e^{r-R}\bigl{(}\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\tfrac{1}{2}\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi\bigr{)}+2\mathop{\rm sinh}\nolimits^{2}R_{+}\sin 2\phi\,,$ (109a)
$(\Delta\hat{Y}_{\rm out})^{2}=4\alpha^{2}\bigl{[}e^{2r}(\mathop{\rm cosh}\nolimits R_{+}\sin\phi+\mathop{\rm sinh}\nolimits R_{+}\sin 2\phi)^{2}+e^{-2R}(\mathop{\rm cosh}\nolimits R_{+}\cos\phi+\mathop{\rm sinh}\nolimits R_{+}\cos 2\phi)^{2}\bigr{]}\\ +4(\mathop{\rm cosh}\nolimits^{2}R_{+}\sin^{2}2\phi+\cos^{2}2\phi)\mathop{\rm sinh}\nolimits^{2}R_{+}\,.$ (109b)
Let $R=0$ and $|\phi|\ll 1$. In this case,
@@
Acknowledgements.
This work of D.S., V.L., B.N., and F.K. was supported by the Theoretical Physics and Mathematics Advancement Foundation “BASIS” Grant #23-1-1-39-1.

### v2

@@ Appendix B Threshold detection, two-arm interferometer case
Let the states $|{\psi_{k_{1},k_{2}}}\rangle$, where $k_{1,2}=1,2,\dots$, complement $|{\psi_{0,0}}\rangle$ to the two-dimensional full orthonormal set:
@@
the operators $\hat{\mathcal{D}}_{1}$, $\hat{\mathcal{S}}_{1,2}$ are given by Eqs. (43) and (44), and $|{n_{1},n_{2}}\rangle$ are the two-dimensional Fock states. Correspondingly, Eq. (87) gives:
@@
$\hat{\mathcal{V}}^{\dagger}\hat{Y}\hat{\mathcal{V}}=\hat{\mathcal{Y}}_{2}+\hat{\mathcal{Y}}_{2}^{\dagger}=-2[\sqrt{2}\alpha\hat{p}_{2}e^{r}+(\hat{x}_{1}\hat{p}_{2}+\hat{x}_{2}\hat{p}_{1})\mathop{\rm sinh}\nolimits R_{+}]$ (94)
(we omit for brevity the additional operator $Q$, see discussion after Eq. (82)).
@@
Let now $R=0$ and $|\phi|\ll 1$. In this case, the cumbersome but straightforward calculations give that